# Simulations of dielectric permittivity of water by Machine Learned Potentials with long-range Coulombic interactions

### v1

@@ V CONCLUSIONS
Using linear response theory, we compute the dielectric constant through both the dipole fluctuation method and the Kirkwood dipole correlation formalism. Three electric boundary conditions—metallic, insulating, and Kirkwood-Fröhlich—are explicitly examined. From the dipole fluctuation method, we extract characteristic polarization relaxation times of approximately 0.3 psps (longitudinal) and 8 psps (transverse). We also introduce a unified approach to determine the Kirkwood correlation length ($\sim$ 16 Å) and Kirkwood correlation factor ($\sim 2.4$) based on the asymptotic behavior of dipole correlation functions across all boundary conditions. Consistently, our simulations yield a dielectric constant of approximately 102 from both computational approaches, irrespective of the boundary condition applied.

### v2

@@ V CONCLUSIONS
Using linear response theory, we compute the dielectric constant through both the dipole fluctuation method and the Kirkwood dipole correlation formalism. Three electric boundary conditions—metallic, insulating, and Kirkwood-Fröhlich—are explicitly examined. From the dipole fluctuation method, we extract characteristic polarization relaxation times of approximately 0.3 $ps$ (longitudinal) and 8 $ps$ (transverse). We also introduce a unified approach to determine the Kirkwood correlation length ($\sim$ 16 Å) and Kirkwood correlation factor ($\sim 2.4$) based on the asymptotic behavior of dipole correlation functions across all boundary conditions. Consistently, our simulations yield a dielectric constant of approximately 102 from both computational approaches, irrespective of the boundary condition applied.